OpenStudy (anonymous):

Solve the following system of equations and show all work. y = x2 + 3 y = x + 5

OpenStudy (anonymous):

I know we have to multiply them by something to be able to cross the variables out or something, I'm not sure

OpenStudy (anonymous):

i have a better idea

OpenStudy (anonymous):

since the y's have to be the same, set \[x^2+3=x+5\] and solve the resulting quadratic equation

OpenStudy (anonymous):

\[x^2-x-2=0\] etc

OpenStudy (anonymous):

x=2, -1?

OpenStudy (anonymous):

seems reasonable find the y values too

OpenStudy (anonymous):

I actually do not now how

OpenStudy (anonymous):

I skimmed over everything and i'm regretting it at the moment

OpenStudy (anonymous):

of course you know how!

OpenStudy (anonymous):

you know \(x=2\) replace \(x\) by 2 in either \(y=x^2+3\) or \(y=x+5\)

OpenStudy (anonymous):

you will get the same number in both cases (since they intersect there) then repeat with \(x=-1\)

OpenStudy (anonymous):

you can do it in your head right?

OpenStudy (anonymous):

Yes, 7.

OpenStudy (anonymous):

so one solution is \((2,7)\) and the other you get when \(x=-1\)

OpenStudy (anonymous):

I'm not fully understanding, I'm sorry :( could you explain?

OpenStudy (anonymous):

you found two solutions for \(x\) right?

OpenStudy (anonymous):

one is \(x=2\) and if \(x=2\) the \(y=7\) in both equations so one point of intersection is \((2,7)\)

OpenStudy (anonymous):

the other solution you got was \(x=-1\) and if you replace \(x\) by \(-1\) in either equation above you get \(4\)

OpenStudy (anonymous):

since \(y=-1+5=4\) and also \(y=(-1)^2+3=4\)

OpenStudy (anonymous):

therefore the other point of intersection (there are two of them, one for each \(x\)) is \((-1,4)\)

OpenStudy (anonymous):

how is that for an explanatation? not sure i can do any better except maybe it is worth mentioning that the graph of \(y=x^2+3\) is a parabola, and it intersects the graph of \(y=x+5\) (a line) in two places

OpenStudy (anonymous):

So the final answer would be (-1,4)?

OpenStudy (anonymous):

no there are two answers

OpenStudy (anonymous):

And (2,7)?

OpenStudy (anonymous):

righhhhhht

OpenStudy (anonymous):

Great. Thank you :)